OpenStudy (anonymous):

A class of 19 pupils has five more girls than boys. Let n = the number of boys. n + (n + 19) = 5 n + (n + 5) = 19 n - (n + 5) = 19 n + (n - 19) = 5

OpenStudy (alesia/bailey):

n+(n+19)=5

OpenStudy (alesia/bailey):

n+(n+19)=5 -19 -19 n+n=-14 2n=-14 2n/2 (crosses out the 2 which just leaves the n) 2n/2=-14/2 n=-7 I don't know if that's right.. but you could mess around with it

OpenStudy (alesia/bailey):

and the rest are the same way. if I did it right..